# Low Temperature Properties of Low-Loss Macroscopic Lithium Niobate Bulk Acoustic Wave Resonators

### v2

@@ VIII Performance at Sub-Kelvin Temperature
The cryostat system as described in sectionII can be further cooled from 4 K down to sub-Kelvin temperatures due to helium dilution. This provides an opportunity to explore the lithium niobate system at extreme cryogenic temperatures approaching the standard quantum limit. Taking a look at figure 6 the quality factors greatly suffer when cooled further to 10 mK. This is due to the increased presence of non linear dissipation channels, such as regimes dominated by self induced absorption. The stronger presence of non linear dissipation is provided by the thermal un-saturation of two level system states due to microsopic impurities. The presence of these states make it difficult to operate the system below nonlinear amplitude thresholds, resulting in general increases to dissipation.
@@ IX Conclusion
Z-cut lithium niobate provides a promising material for macroscopic BAW resonators designed with the intention of testing areas of fundamental physics, demonstrating large quality factors for bulk acoustic modes. However, the existence of microscopic impurities sites in the crystalline bulk puts limitations on the ultimate performance for high overtones. Additionally, the strong thermoelastic damping of the material results in further attenuation at low frequencies in comparison to crystalline quartz. We demonstrate that the losses in our system are well understood, with intrinsic quality factors adhering to models predicted by theory. Future improvements to lithium niobate resonators can be made by reducing atomic impurity levels. Treatments such as DC sweeping, and improvements to crystal growth procedures have been successfully to produce higher quality material in other systems. Pending such engineering improvements lithium niobate BAW resonators will see high overtone quality factors tend towards the billions, reaching the intrinsic limit set by thermoelastic damping and phonon-phonon scattering. Such high quality factors will be beneficial to various fundamental physics experiments hoping to exploit the large mass macroscopic phonon modes provided by this system.

### v3

@@ VIII Performance at Sub-Kelvin Temperature
The cryostat system as described in section II can be further cooled from 4 K down to sub-Kelvin temperatures due to helium dilution. This provides an opportunity to explore the lithium niobate system at extreme cryogenic temperatures leading to the quantum regime where the thermal occupation number approaches unity. Taking a look at figure 7 the quality factors greatly suffer when cooled further to 10 mK. Some reduction in $Q$ can occasionally be compensated by driving to larger amplitudes. This feature is reminiscent of the saturation of TLS dissipation, however the effect observed here does not follow well known TLS models [31]. Instead, the extra dissipation at this temperature could be described by the oscillator state being deeper into the non-linear regime of section VII and thus the saturation comes as the system enters the regime of self-induced transparency. We observe that the critical applied drive power for self-induced absorption to be lower at 10 mK than at 4 K for nearly all modes.
At sub-Kelvin temperatures a reduction of phonon-phonon losses causes an increase to the intrinsic energy stored in the crystal, and thus the acoustic motion enters a deeper non-linear regime when compared to 4 Kelvin temperatures for the same incident drive power. Such phenomena is well known when it comes to the duffing effect in quartz [56, 59], and could be the cause for the stronger non-linear dissipation observed here. It is important to note that in these strong non-linear regimes the quality factor becomes harder to strictly define with multiple channels contributing to intrinsic dissipation.
@@ IX Conclusion
Z-cut lithium niobate provides a promising material for macroscopic BAW resonators designed with the intention of testing areas of fundamental physics, demonstrating large quality factors for bulk acoustic modes. However, the existence of microscopic impurities sites in the crystalline bulk puts limitations on the ultimate performance for high overtones. Additionally, the strong thermoelastic damping of the material results in further attenuation at low frequencies in comparison to crystalline quartz. We demonstrate that the losses in our system are well understood, with intrinsic quality factors adhering to models predicted by theory. Future improvements to the quality factors of modes $f_{r}>30$ MHz in lithium niobate resonators can be made by reducing atomic impurity levels. Treatments such as DC sweeping, and improvements to crystal growth procedures have been successfully to produce higher quality material in other systems. Pending such engineering improvements lithium niobate BAW resonators will see high overtone quality factors tend upwards of $10^{7}$, reaching the intrinsic limit set by thermoelastic damping. Such high quality factors will be beneficial to various fundamental physics experiments hoping to exploit the large mass macroscopic phonon modes provided by this system.
X Acknowledgments
This research was supported by the ARC Centre of Excellence for Engineered Quantum Systems (EQUS, No.CE170100009) along with support from the Defence Science and Technology Group (DSTG) as part of the EQUS Quantum Clock Flagship program. Additional support provided by the ARC Centre of Excellence for Dark Matter Particle Physics (CDM, No. CE200100008).